OpenStudy (anonymous):

17 = a -3 Choose all that are correct. -14 14 -20 20

OpenStudy (anonymous):

Would it be 20 and 14?

OpenStudy (anonymous):

just 20

OpenStudy (anonymous):

there is only 1 answer to this equation

OpenStudy (anonymous):

14-3=11, not 20

OpenStudy (anonymous):

* 17

OpenStudy (anonymous):

17 = a - 3 17+3 = a 20 = a

OpenStudy (anonymous):

Thank you